Please show work.
Consider the function
f(x)=x^5-x^3
a Determine the behaviors as x→∞ and as x→-∞
b Find all zeros. What is the multiplicity of roots (zeros)?
c When will the function cross the x-axis? When will it cross the y-axis?

Answer by ewatrrr(24785)


Hi
f(x)=x^5-x^3 = x^3(x^2 - 1) = x^3(x-1)(x+1)